OpenStudy (anonymous):

solving by addition: find X and Y 2x+3y=7-2x 7x+2y=9

OpenStudy (anonymous):

2x+3y=7-2x 4x+3y=7

OpenStudy (anonymous):

7x+2y=9 4x+3y=7

OpenStudy (anonymous):

x = 1, y = 1

OpenStudy (anonymous):

starting with the second system by Libniz, you could multi. the first equation by 3, and the second equation by -2; that way the y coefficients will be opposites and will be eliminated when the two equations are added together: 21x+6y=27 -8x-6y=-14 ======== 13x =13 x=1 now back sub to get y; i'll choose the original second equation (you know that you can use any equation that has a y in it) 7(1)+2y=9 (sub in 1 for x) 2y=9-7 (subtract 7 from both sides) y=1 (divide through by 2) the answer may be written as x=1, y=1; some like it as (1,1)